26.

An investment of $15,000 is made for three years at 4.5%

compounded quarterly.


a)

Determine the maturity value of the investment.

b)

What is the amount of compound interest earned on the

investment?


1
Expert's answer
2021-08-02T06:06:11-0400

"FV=PV(1+\frac{r}{n})^{n×t}"


"=15000(1+\frac{0.045}{4})^{4×3}"


"=15000(1.01125)^{12}"


"=15000(1.14367)=\$17155.05"

Maturity value of the investment"=\$17155.05"


b) The amount of compound interest earned on the investment

"=\$17155.05-\$15000=\$2155.05"